A car is driven off a cliff at a speed of 50km an hour. It lands .350 km away from the cliff. How tall is the cliff?
*I usually solve these problems knowing the height of the cliff and asked to find the speed of the car, now I'm confused.
thanks!

Answer by ankor@dixie-net.com(22740)

A car is driven off a cliff at a speed of 50km an hour.
It lands .350 km away from the cliff.
How tall is the cliff?
:
Find out how long for a car to travel .350 km at a speed of 50km/hr (seconds)
t = .350/50 * 3600 = 25.2 sec
:
How far will an object fall in 25.2 sec (h = height of the cliff)
h = 4.9(25.2^2)
h = 4.9 * 635
h = 3111.7 meters
:
:
Seem awful high, but I think that is right!